OpenStudy (anonymous):

The sum of three consecutive integers is 42. Find the Integers.

OpenStudy (anonymous):

x+(x+1)+(x+2)=42??????

OpenStudy (anonymous):

does that look right to anyone?

OpenStudy (isaiah.feynman):

Yes that is absolutely right!!!

OpenStudy (anonymous):

it's nicer IMHO to label the middle \(n\) and so that our three are \(n-1,n,n+1\) -- when you sum them they cancel out nicely

OpenStudy (anonymous):

Isaiah will you please help me with this?

OpenStudy (anonymous):

ok so its x+(x+1)+(x+2)=42

OpenStudy (isaiah.feynman):

Yes, I've solved and verified. You are right!!! Simply solve for x

OpenStudy (anonymous):

I'm trying to remember how to do this

OpenStudy (anonymous):

hmmm

OpenStudy (isaiah.feynman):

When you sum them up it becomes 3x+3=42. Solve for x...

OpenStudy (anonymous):

3x+3=42??

OpenStudy (anonymous):

ok wow that took forever

OpenStudy (isaiah.feynman):

lol

OpenStudy (anonymous):

x=13 got it thanks

OpenStudy (anonymous):

If I didn't need to pass this class I would of left that problem a long time ago

OpenStudy (anonymous):

its not adding up to 42

OpenStudy (isaiah.feynman):

Try again :)

OpenStudy (anonymous):

alternatively, we have \(n-1,n,n+1\) our three consecutive integers and their sum is:$$(n-1)+n+(n+1)=42\\3n=42\\n=\frac{42}3=\frac{2\times21}3=2\times7=14$$hence our integers are \(13,14,15\). they do indeed sum to \(42\)

OpenStudy (anonymous):

they dont add up to 42?

OpenStudy (anonymous):

yeah they do https://www.google.com/search?q=13%2B14%2B15

OpenStudy (isaiah.feynman):

@Algebra2013 13+(13+1)+(13+2) = what??????

OpenStudy (anonymous):

wait a minute here

OpenStudy (anonymous):

I put in 13+(14+1)+(15+2)=42

OpenStudy (anonymous):

I'm an idiot. thanks guys

OpenStudy (anonymous):

hehe we all make silly little mistakes :-p